Symon Meats is looking at a new sausage system with an installed cost of $510,000. This cost will be depreciated straight-line to zero over the project’s five-year life, at the end of which the sausage system can be scrapped for $76,000. The sausage system will save the firm $190,000 per year in pretax operating costs, and the system requires an initial investment in net working capital of $35,000. If the tax rate is 34 percent and the discount rate is 10 percent, what is the NPV of this project? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

rate positively ..

computation of NPV
Year 0 1 2 3 4 5
i cost of system -510000
ii working capital -35000 35000
Operating cash flow
a Saving in cost 190000 190000 190000 190000 190000
b Depreciation 102000 102000 102000 102000 102000
c=a-b Profit before tax 88000 88000 88000 88000 88000
d=c*34% Tax @ 34% 29920 29920 29920 29920 29920
e=c-d Profit after tax 58080 58080 58080 58080 58080
iii=e+b Operating cash flow 160080 160080 160080 160080 160080
iv Post tax salvage value
76000*(1-34%) 50160
v=i+ii+iii+iv Net cash flow -545000 160080 160080 160080 160080 245240
vi PVIF @ 10%             1.0000          0.9091          0.8264          0.7513          0.6830          0.6209
vii Present value     (545,000.00) 145,527.27 132,297.52 120,270.47 109,336.79 152,274.75 114,706.81
therefore NPV =      114,706.81